Date:
Monday, September 16, 2019
Telescope:
CTIO 1.5-m
Submitter:
Roberto Aviles
Submitter Email:
Observer Support:
None
Night Assistant:
None
Problems:
23:27 UT requested move to V3890 Sgr: Tel stopped moving + dome disabled.
Program:
Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
10.5
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
3.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
13.5
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2019-09-16 | 10.5 | 0.0 | 0.0 | 3.0 | 0.0 | 13.5 |
04:50 UT closed: condensation and 77% RH: foggy (CTIO weather station down at 6:30UT.) Reopened at 07:45 UT based on visual inspection + LCO and MEarth stations.
Underline than the seeing was greater than 3 arc-sec, reached at less 7 arc sec.
Thanks to the Mechanics team for their works with the dome shutter.